find the following. assume that variables can represent any real number

square root (a+5)^2

Answer by stanbon(75887)

find the following. assume that variables can represent any real number
square root (a+5)^2
= (a+5)(a+5)
---
= a(a+5) + 5(a+5)
---
= a^2 + 5z + 5a + 25
---
= a^2 + 10a + 25
====================
Cheers,
Stan H.
====================